# Eye Gaze as a Signal for Conveying User Attention in Contextual AI Systems

### v1

@@ 4.2. Contextualized vision-language model queries
Figure 4. Experiments where prior gaze fixation contents are supplied to a VLM along with egocentric images. When many fixations are considered, the model synthesizes image / gaze context to outperform a greedy baseline, which only considers contents from the prompt. Error surfaces in light blue represent 95% confidence intervals.
We constrain the VLM model to respond with a selection from the list of all currently visible objects, making these experiments classification tasks where accuracy = correct selections / all trials. Cases where the VLM response failed to return parseable JSON ($<$1% of trials) were discarded.
We also construct a number of baselines for comparison against the VLM selection accuracy. These baselines implement simple heuristics which select one object out of a number of visible objects and/or fixated objects. The lowest performing baseline is random guessing among all visible objects. We also model random guesses from the list of previously fixated objects, and a greedy strategy which always chooses the most fixated object from a sequence of previously fixated objects. With a single prior fixation, our random guessing baseline instead represents always predicting the subject of the prior fixation.
@@ 4.2.1. E1: “What am I looking at?”
When querying the VLM without supplying any additional context, the model successfully predicts the current fixated object 10.3% (95% CI = [8.3%, 12.3%]) of the time (see Figure 4 (left)). While this surpasses random guessing from visible objects, it is quite low. The model does not surpass a greedy baseline which always returns the immediately preceding fixation (left tail of Random (prior fixations) in Figure 4). VLMs are not expected to excel at this task; they are known to misinterpret context and are prone to multimodal hallucination (Cui et al., 2023; Leng et al., 2024); this task’s specificity is better accomplished from saliency prediction (Kroner et al., 2020; Wang et al., 2024). However, including context from prior gaze greatly improves the model’s ability to predict the current fixation, with a peak accuracy of 24.8% (CI = [22.1%, 27.7%]) at 6 prior fixations. Context-based baselines slightly outperform the VLM with one or few prior fixations, reinforcing that current gaze is contingent on scanpath history (Burlingham et al., 2024a, b). With more context (6+ fixations), the model outperforms baselines, indicating that prior context and image contents are being synthesized, and the combination of contextual cues increase the model’s performance.
@@ 4.2.2. E2: “What am I going to interact with?”
E2 sees similar trends to E1; however, the more contextually-grounded task to predict the object of physical interaction sees greater benefit from the inclusion of ET context. Clearly, prior eye gaze is a strong indicator for interaction, and by supplying preceding gaze context, we can greatly improve the VLM’s ability to understand the user’s actions. We see a peak accuracy of 49.5% (CI = [43%, 56.1%]). As evident by this and the stronger baseline performances, gaze is tightly coupled with the onset of interaction. Note that we query only at positive examples where an interaction does take place, and the inclusion of a null case could have led to the model raising false positives / negatives.
@@ 5. Discussion
In this work, we explored the future role of eye tracking in wearable contextual AI systems: as the main signal to convey user attention relative to the physical world. Analyzing household tasks in a digitally twinned environment, we benchmarked the ET signal quality requirements for a contextual AI system to accurately and reliably sense objects. We then referenced gaze data alongside object labels to convey a user’s scanpath history to a VLM, prototyping ET’s role in aggregating contextual cues. Our experiments show that prior gaze fixations (scanpath history) enhance the VLM’s understanding of image contents, and are a strong prior signal for human object interaction.
Contextual cues from ET provide clear benefits, as evident in subsection 4.2. We expect ET’s value to become more prominent in future models which are trained specifically for egocentric understanding and / or with eye gaze as a direct input (Koorathota et al., 2023). Our findings build on prior works (Toyama et al., 2012; Burlingham et al., 2024b), making it evident that human actions and gaze patterns display temporal dependencies contingent on prior gaze / actions, similar to the dependencies in written language. If we can effectively convey previous context, VLMs may become able to better infer current / future context based on the patterns present in prior behavior.
In our ET signal quality benchmark, we found that nearby objects tend to be within the visual size requirements for current wearable ET devices to effectively sense them. Yet, current systems may struggle to detect edge cases (physically small or far away objects), as shown by the tail end of our distributions (95% coverage of fixated and mid-field objects being $\sim$1.15°$\degree$ for radius and $\sim$0.66°$\degree$ for minor axis). In the future, it could be possible to supplement ET accuracy in the future, possibly with contextual cues (Bi and Zhai, 2013) for error correction, additional sensors (Wei et al., 2023), or abstracting eye gaze alongside language (team et al., 2024) to improve wearable ET for contextual AI applications.
@@ 5.1. Conclusion
Eye tracking improves agent understanding of the physical world and aligns this understanding with users. Our results suggest that for close by scenarios, such as active grabbing / touching of objects and gaze selection, current ET systems can consistently place fixations on objects and convey relevant information to VLM agents. We showcased direct benefits when supplying scanpath history, greatly improving VLM understanding of the physical space when using gaze-contingent context from the nearby past. With ET context, future contextual agents will have a greater understanding about the world and about the users’ attentive state and intents.

### v2

@@ 4.2. Contextualized vision-language model queries
Figure 4. Experiments where prior gaze fixation contents are supplied to a VLM along with egocentric images. When many fixations given as context, the model can synthesize information between the image + gaze context to outperform a greedy baseline that only considers contents from the prompt. The error surfaces in light blue represent 95% confidence intervals.
Because we constrain the VLM model to respond by selecting from the list of all currently visible objects, these experiments are classification tasks where accuracy = correct selections / all trials. Cases where the VLM response failed to return parseable JSON ($<$1% of trials) were discarded.
We compare against a number of baselines to see if the model effectively uses the context supplied. These baselines implement simple heuristics on the image and gaze context. The lowest performing baseline is random guessing among all visible objects. We also implement random guesses from the list of previously fixated objects, and a greedy strategy which always chooses the most fixated object from a sequence of previously fixated objects. Note these baselines still use context provided from ET, but make no attempt to synthesize with the egocentric image for better world understanding.
@@ 4.2.1. E1: “What am I looking at?”
When querying the VLM only with the egocentric image as context, the model successfully predicts the current fixated object 10.3% (95% CI = [8.3%, 12.3%]) of the time (see Figure 4 (left)). While this surpasses random guessing from visible objects, it is quite low. An effective strategy was to always return the immediately preceding fixation (left tail of Random (prior fixations) in Figure 4). VLMs are not expected to excel at gaze prediction, as they are known to misinterpret context or hallucinate (Cui et al., 2023; Leng et al., 2024). However, including context from prior gaze greatly improves the model’s ability to predict the current fixation, with a peak accuracy of 24.8% (CI = [22.1%, 27.7%]) at 6 prior fixations. The gaze context-based baselines slightly outperform the VLM with one or few prior fixations, reinforcing that current gaze is contingent on scanpath history (Burlingham et al., 2024a, b). With more context (6+ fixations), the model begins to outperforms baselines, indicating that prior context and image contents are being synthesized, and the combination of contextual cues increase the model’s performance.
@@ 4.2.2. E2: “What am I going to interact with?”
E2 sees similar trends to E1; however, the more contextually-grounded task to predict the object of physical interaction sees greater benefit when including ET context. Clearly, prior eye gaze is a strong indicator for interaction, and by supplying historical gaze context, we can greatly improve the VLM’s ability to understand the user’s actions. We see a peak accuracy of 49.5% (CI = [43%, 56.1%]). As evident by this and the stronger baseline performances, gaze is tightly coupled with the onset of interaction. Note that we query only at positive examples where an interaction does take place, and the inclusion of a null case could have led to the model raising false positives / negatives.
@@ 5. Discussion
We expect that ET’s value would become even more prominent in future models which are trained specifically for egocentric understanding and / or with eye gaze as a direct input (Koorathota et al., 2023). Our findings, building on prior works (Toyama et al., 2012; Burlingham et al., 2024b), evidence that human actions and gaze patterns display temporal dependencies contingent with previous actions, similar to the dependencies in written language. If we can effectively convey the traces of human attention and actions, VLMs may become able to better infer current / future context based on the patterns present in prior behavior.
Our ET signal quality benchmark measures the likelihood of sensing objects, but has little considerations of edge cases (such as very small or very far objects). In the future, supplementary computations might aid the ability to place gaze on the correct object, possibly via contextual cues (Bi and Zhai, 2013) for error correction or additional sensors (Wei et al., 2023). These could alleviate ET sensor quality from becoming a bottleneck when using gaze to model human attention, specifically in more challenging contexts such as outdoors, where objects tend to be much further.
@@ 5.1. Conclusion
We investigated eye tracking signals’ ability to improve multimodal agents’ understanding of the physical world. Our results suggest that for close by scenarios, such as active grabbing / touching of objects and gaze selection, current ET systems could consistently place fixations on objects and convey relevant information to VLM agents. In our experiments, we saw direct benefits when adding scanpath history to queries. Given these findings, future contextual agents which receive signals about user attentive state may obtain a greater understanding about the world, and better align with user intent, improving the usability of such systems.